OpenStudy (anonymous):

find the indefinite integral x(2x+1) dx

OpenStudy (unklerhaukus):

\[\int x(2x+1)\text dx\]?

OpenStudy (lgbasallote):

distribute first

OpenStudy (unklerhaukus):

\[=\int(2x^2+x)\text dx\] \[=...\]

OpenStudy (lgbasallote):

it should be an easy integral then

OpenStudy (lgbasallote):

so getting any ideas @Lachlan1996 ?

OpenStudy (unklerhaukus):

\[=\frac{2x^{2+1}}{2+1}+\frac{x^{1+1}}{1+1}+c\]\[=\dots\]

OpenStudy (lgbasallote):

is that your final answer @UnkleRhaukus /:)

OpenStudy (lgbasallote):

^by that i mean are you sure about that...

OpenStudy (unklerhaukus):

um?

OpenStudy (unklerhaukus):

, yes/?

OpenStudy (anonymous):

oh my god im so sorry, i had my laptop closed and i had moved on from this question. My apologies everyone and thankyou for your assistance. @UnkleRhaukus

OpenStudy (anonymous):

and @igbasallotte

OpenStudy (anonymous):

i mean @lgbasallote